OpenStudy (anonymous):

A starburst has the following equations: y=2x y=x y= (1/2)x y=0 y=(-1/2)x y=-x y=-2x x=0 and all of these lines intersect at the orgin (0,0). If one was to move the starburst so they all intersect at the point (4,2), how would they do so and what would the new euqations be?

OpenStudy (anonymous):

equations: y=2x. y=x. y=(1/2)x. y=0. y=(-1/2). y=-x. y=-2x. x=0.

OpenStudy (anonymous):

y=2(x-4)+2 y=(x-4)+2 y=1/2(x-4)+2 y=2 y=-1/2(x-4)+2 y=-(x-4)+2 y=-2(x-4)+2 x=4 Basically to move every equation along the x-axis you have to put (x-n) to move it to the right and (x+n) to move it to the left. Then all you have to do is add +n to move the y-axis up and -n to move the y-axis down. So the original formula is y=mx+b, if you make the desired intersection point (a,b), then, y=m(x-a)+b. Since y=2 is a horizontal line because the slope is 0 and x=0 so therefore has not slope; x=4 is a vertical line because it is undefined on the y-axis.